Simplify the expression to lowest terms. 20x^4y^6/50x^6y^9 ( this is supposed to be a fraction, the only way I write it.

Answer by stanbon(75887)

Simplify the expression to lowest terms. 20x^4y^6/50x^6y^9 ( this is suppose to be a fraction, the only way I write it.
------------------
20x^4y^6/50x^6y^9
---
= (20/50)(x^4/x^6)(y^6/y^9)
----
= (2/5)(1/x^2)(1/y^3)
---
= 2/(5x^2y^3)
======================
Cheers,
Stan H.
----------